You have $6,000 to invest into two stock funds. The first fund pays 5% annual interest and the second account pays 9% annual interest. If after a year you have made $380 in interest, how much money did you invest in each account?
Answer by Boreal(15235)

x in the first
6000-x in the second
.05x+.09(6000-x)=380
.05x+540-.09x=380
-0.04x=-160
x=$4000 @ 5% is $200 ANSWER
6000-x=$2000 @9%=$180 ANSWER